An oxide of nitrogen contains 63.1% oxygen and has a molar mass of 76.0 g/mol. What is the empirical

formula for this compound?



1
Expert's answer
2021-05-14T03:11:58-0400

Take a hypothetical 100-gram sample of the gas and convert each fractional weight to moles.


"\dfrac{63.1g}{15.9994(\frac{g}{mole})}=3.9438979 \;moles\;of\;O"


"\dfrac{36.9\;g}{14.0067(\frac{g}{mole})}=2.6344525\;moles\;ofN"


Divide by the smaller number of moles:


"O:\dfrac{3.9438979}{2.63445351}=1.49704593"


"N:\dfrac{2.63445351}{2.63445351}=1"


Multiply by 2 to make integer ratios:

N2O3 This is the empirical formula.



Answer:

N2O3 has a molar mass of 76.0117 "(\frac{g}{mole})", so it is molecular formula too.